OpenStudy (anonymous):

(x + 8)/(x +6)(x -8) * 4x(x + 10)/( x + 10)

OpenStudy (luigi0210):

\[\frac{ x+8 }{ (x+6)(x-8) }*\frac{ 4x(x+10) }{ x+10) }\] This?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

What are we doing exactly? simplifying? multiplying?

OpenStudy (anonymous):

We doing both

OpenStudy (luigi0210):

Well first simplify the equation.

OpenStudy (anonymous):

Okay I really don't know how to simplify this one

OpenStudy (luigi0210):

Well look at the equations, anything look the same on the top/bottom?

OpenStudy (anonymous):

yea the number on the right (x+10) (x+10)

OpenStudy (luigi0210):

Good, now cancel them out and see what the leaves you with

OpenStudy (anonymous):

it leaves you with \[(x + 8)/(x +6)(x -8) * 4x\]

OpenStudy (luigi0210):

correct! Now just multiply it out

OpenStudy (anonymous):

\[x ^{3}-384*4x\]

OpenStudy (luigi0210):

:l

OpenStudy (anonymous):

that's not the right answer

OpenStudy (luigi0210):

Nope ^_^'

OpenStudy (anonymous):

what you suppose to get when you multiply it.

OpenStudy (luigi0210):

\[\frac{ x+8 }{ (x+6)(x-8) }*4x=\frac{ 4x(x+8) }{ x^2-2x-48 }\]

OpenStudy (luigi0210):

Finish it by multiplying the numerator

OpenStudy (anonymous):

\[\frac{ 4x ^{2} + 32x }{ (x+6) (x-8) }\]

OpenStudy (luigi0210):

Yup :)

OpenStudy (anonymous):

that's the answer

OpenStudy (luigi0210):

Yes sir!

OpenStudy (anonymous):

thank you soo much

OpenStudy (luigi0210):

You are very welcome :)